OpenStudy (anonymous):

please help i dont understand how to solve this ^3√32 please help i dont understand how to solve this ^3√32 @Mathematics

OpenStudy (phi):

I assume you mean \[\sqrt[3]{32}\] to do these, break up the 32 into its factors. In this case it is 2^5 or \[2\cdot 2\cdot 2\cdot 2\cdot 2= (2\cdot 2\cdot 2)(2\cdot 2)\] try to group sets of three. Here we have a triple of 2's. We can take those out of the cube root (and reduce the three 2's down to 1 2). We get \[2\sqrt[3]{4}\]